# Do language models accommodate their users? A study of linguistic convergence

### v1

@@ Abstract
While large language models (LLMs) are generally considered proficient in generating language, how similar their language usage is to that of humans remains understudied. In this paper, we test whether models exhibit linguistic convergence, a core pragmatic element of human language communication, asking: do models adapt, or converge, to the linguistic patterns of their user? To answer this, we systematically compare model completions of exisiting dialogues to the original human responses across sixteen language models, three dialogue corpora, and a variety of stylometric features. We find that models strongly converge to the conversation’s style, often significantly overfitting relative to the human baseline. While convergence patterns are often feature-specific, we observe consistent shifts in convergence across modeling settings, with instruction-tuned and larger models converging less than their pretrained counterparts. Given the differences between human and model convergence patterns, we hypothesize that the underlying mechanisms for these behaviors are very different.
@@
Terra Blevins1,3     Susanne Schmalwieser2     Benjamin Roth1 1Faculty of Computer Science, University of Vienna 2Institute of German Studies, University of Vienna 3Khoury College of Computer Sciences, Northeastern University Correspondence: terra.blevins@univie.ac.at
@@ 1 Introduction
In this work, we investigate this through the lens of accommodation, the process by which people adjust their speech or writing style based on the identity of their interlocutor Giles et al. (1991). While accommodation takes many forms, in this paper we focus on linguistic convergence, or how similar the target’s language is to their interlocutor’s Niederhoffer and Pennebaker (2002). As the most prevalent use case for LLMs is via a chat agent, convergence acts as an interesting case study into the novel setting of human-model interaction. Thus, we seek to answer the following questions: Do LLM-based chatbots adapt their language use to mirror that of their users? And does this behavior mirror trends observed in human accommodation or follow different patterns?
@@
To study convergence in LLMs, we perform a systematic analysis of model-generated responses to human-authored conversations, which we collect by prompting various LLMs to complete turns in existing conversations drawn from dialogue corpora. We then analyze these responses for a series of stylometric features used to study accommodation in human linguistics, in order to estimate the degree of convergence LLMs exhibit towards their context relative to randomly-sampled utterances and a human control (i.e., the gold text in turns that the model replaces).
@@
Our experiments show that models strongly adapt their outputs’ style to their context. We also find differences in convergence across model size and training schemes, with larger and instruction-tuned models converging less than smaller or solely pretrained models. Given that pretrained models often overconverge on these linguistic markers relative to the human baseline, this indicates that instruction-tuned models are frequently more in line with human behavior. Furthermore, these convergence patterns are feature-specific, with models adapting to their context to different extents across metrics. While this varied behavior across features is also common in human accommodation, the variance we observe in the models often differs from that of the human baseline, indicating subtler differences in human and model convergence. We then conclude by discussing the implications of this work towards understanding the linguistics of LLMs.
@@
To summarize, this paper presents the following primary contributions: (1) we develop a novel, synthetic paradigm for testing model behavior with respect to a paired human control, which we apply to studying linguistic convergence in LLMs; and (2) we present a quantitative analysis of into how language models converge to the style of their context, with insights into how modeling factors affect this phenomenon and the implications it has for interpreting model-generated text. Our findings demonstrate how models do and do not adapt to their inputs across various stylistic features, providing broader insights into the relationship between human and model language use. [^1] The source code and model generations can be found at https://github.com/blvns/llm-convergence.
@@ 3 Measuring Linguistic Convergence in Language Models
In addition to directly measuring model convergence, we also compare the models’ behavior to two baselines: the human baseline, which considers how well the original utterance that $r_{t}$ replaces accommodates the prior text, [^2] I.e., establishing the accommodation exhibited by humans in the same setting. and the random baseline, which calculates how much a random utterance drawn from a different conversation in the dataset accommodates $r_{t-1}$.
@@ 3.1 Linguistic Indicators of Convergence
We measure the following features to characterize whether models alter their linguistic style to match that of their users, drawing on both human accommodation research and other stylometric features:
@@
LIWC Agreement A standard measure of linguistic accommodation is the frequency of LIWC Chung and Pennebaker (2012) function word classes (e.g. Danescu-Niculescu-Mizil and Lee, 2011). Here, we consider the LIWC2007 classes considered in Ireland et al. (2011) (personal and impersonal pronouns, articles, conjunctions, prepositions, auxiliary verbs, frequently used adverbs, negations, and quantifiers), also using the $LSM$ metric to calculate how well each LIWC category distribution in generated responses aligns with prior turns. We both report the mean LIWC agreement across categories and provide a fine-grained analysis of each category in Section 5.3.
PROPN Overlap We calculate the overlap (percentage) of proper nouns between the text generated by the model and the preceding turn. We expect that language users converging more with their interlocutor will have a higher overlap.
Token Novelty We evaluate the percentage of tokens novel relative to the reference utterance, where a smaller percentage of novel tokens would indicate that the model is adapting more to the user interacting with it. This is measured as $|w_{t}\cap w_{t-1}|/|w_{t}|$, where $w_{x}=\{w\in r_{x}\}$.
Utterance length, LIWC Agreement, and PROPN Overlap are symmetric metrics and don’t account for directional influence. Given the nature of our experiments (the prior context outside of generations is fixed and can not be affected by the model), we consider this a reasonable assumption to hold.
@@ 4 Experimental Setup
Datasets We perform our dialogue prompting experiments on three English datasets: DailyDialog Li et al. (2017), containing conversations about daily life as written by English language learners; NPR Majumder et al. (2020), a dataset of radio interview transcripts; and the Movie Corpus Danescu-Niculescu-Mizil and Lee (2011), which contains (fictional) conversations scraped from movie scripts.
@@ 5.3 Fine-grained Analysis of LIWC Categories
Table 2: Examples of model responses: pretrained models overconverge on quantifiers (Case A) and instruction-tuned models overuse (and thus underconverge on) conjunctions (Case B).
@@ 5.4 Stepwise Analysis of Linguistic Convergence
We find that human convergence scores fluctuate across timesteps; this is unsurprising, as alternating timesteps are uttered by the same speaker $S_{y}$ as the considered utterance $r_{t}$, indicating that each speaker’s linguistic patterns agree more with themselves than with their interlocutors. Interestingly, we see similar patterns across time with the language models for most metrics. This finding suggests that not only do language models adapt to their context, they also differentiate this adaptation to different roles within the dialogue.
@@ 6 Discussion
Throughout our analysis, we compare model convergence to that of the human baselines to contextualize the models’ behavior. However, while in some cases the model converges to its context similarly to human utterances, we emphasize that the observed similarities do not necessarily indicate that the underlying causes of these behaviors are the same. Human accommodation is driven by speakers (often unconsciously) altering their speech to foster social and communicative goals Giles et al. (1991). In contrast, language models do not have these same underlying communication goals when generating text.
@@
Instead, an important consideration is how the user will interpret texts from models that appear to accommodate them as a human interlocutor would, such as in the case of instruction-tuned models. Appropriate model style has been shown to facilitate successful chatbot interactions Chaves et al. (2019); Thomas et al. (2020), and Bhatt and Rios (2021) finds that users tend to accommodate models more when they successfully produce topically relevant outputs, treating them more like human conversation partners. Thus, model convergence that is more in line with humans (rather than sycophantic, as overconvergence may appear) will likely lead to higher trust in the model, even if the true capabilities of the model are unreliable.
@@ 7 Conclusion
Thus, we consider the extent to which model convergence is even related to human accommodation. Finer-grained analysis on LIWC categories (a prototypical feature for studying accommodation) shows that models exhibit very different patterns on these features than humans, suggesting that the underlying mechanisms for these behaviors are likely very different. Therefore, while this work primarily considers human behavior as a reference for characterizing the models’ generations, future work should characterize these differences further and examine the underlying causes of the observed convergence in LMs.
@@
We approximate the participation of LLMs in user-driven dialogues by having them complete turns in existing dialogue datasets. While this approach has some experimental advantages (i.e., we can directly compare human accommodation features with the model’s responses in the same context), it also presents some limitations. Specifically, the model’s responses may be biased due to participating only in the later portion of (fixed) conversations. Future work should supplement our experimental setting with more focused user studies to validate whether our findings hold in these cases.
While we perform experiments on sixteen language models and three dialogue datasets, it remains an open question how larger models (>>70B parameters) and models post-trained on other objectives adapt to their users. The considered datasets cover various styles of conversations, but these differ from how some users interact with the models (i.e., information-seeking). Thus, future work should confirm whether these findings hold up in interactive user studies. Finally, while we test for an array of stylometric features, it is possible that the model’s behavior of other aspects of style may differ.

### v2

@@ Abstract
While large language models (LLMs) are generally considered proficient in generating language, how similar their language usage is to that of humans remains understudied. In this paper, we test whether models exhibit linguistic convergence, a core pragmatic element of human language communication: do models adapt, or converge, to the linguistic patterns of their user? To answer this, we systematically compare model completions of existing dialogues to original human responses across sixteen language models, three dialogue corpora, and various stylometric features. We find that models strongly converge to the conversation’s style, often significantly overfitting relative to the human baseline. While convergence patterns are often feature-specific, we observe consistent shifts in convergence across modeling settings, with instruction-tuned and larger models converging less than their pretrained and smaller counterparts. Given the differences in human and model convergence patterns, we hypothesize that the underlying mechanisms driving these behaviors are very different.
@@
Terra Blevins1,2     Susanne Schmalwieser3     Benjamin Roth2,3 1Khoury College of Computer Sciences, Northeastern University 2Faculty of Computer Science, University of Vienna, Vienna, Austria 3Faculty of Philological and Cultural Studies, University of Vienna, Vienna, Austria Correspondence: t.blevins@northeastern.edu
@@ 1 Introduction
In this work, we investigate this through the lens of accommodation, the process by which people adjust their speech or writing style based on the identity of their interlocutor Giles et al. (1991). While accommodation takes many forms, in this paper we focus on linguistic convergence, or how similar the target’s language is to that of their interlocutor Niederhoffer and Pennebaker (2002). As the most prevalent use case for LLMs is via a chat agent, convergence acts as an interesting case study into the novel setting of human-model interaction. Thus, we seek to answer the following questions: Do LLM-based chatbots adapt their language use to mirror that of their users? And does this behavior mirror trends observed in human accommodation or follow different patterns?
@@
To study convergence in LLMs, we perform a systematic analysis of model-generated responses to human-authored conversations, collected by prompting various LLMs to complete turns in existing conversations drawn from dialogue corpora. We then analyze these responses for a series of stylometric features used to study accommodation in human linguistics, in order to estimate the degree of convergence LLMs exhibit towards their context relative to randomly sampled utterances and a human control (i.e., the gold text in turns that the model replaces).
@@
Our experiments show that models strongly adapt their output style to their context. We also find differences in convergence across model size and training schemes, with larger and instruction-tuned models converging less than smaller or solely pretrained models. Given that pretrained models often overconverge on these linguistic markers relative to the human baseline, this suggests that instruction-tuned model responses are often more in line with human behavior. Furthermore, these convergence patterns are feature-specific, with models adapting to their contexts to varying degrees across metrics. While this varied behavior across features is also common in human accommodation, the variance we observe in the models often differs from that of the human baseline, indicating subtler differences in human and model convergence. We then conclude by discussing the implications of this work towards understanding the linguistics of LLMs.
@@
To summarize, this paper presents the following primary contributions: (1) we develop a novel, synthetic paradigm for testing model behavior with respect to a paired human control, which we apply to studying linguistic convergence in LLMs; and (2) we present a quantitative analysis of how language models converge to the style of their context, with insights into how modeling factors affect this phenomenon and the implications it has for interpreting model-generated text. Our findings demonstrate how models do and do not adapt to their inputs across various stylistic features, providing broader insights into the relationship between human and model language use. [^1] The source code and model generations can be found at https://github.com/blvns/llm-convergence.
@@ 3 Measuring Linguistic Convergence in Language Models
Baselines In addition to directly measuring model convergence, we also compare the models’ behavior to two baselines: the human baseline, which considers how well the original utterance that $r_{t}$ replaces accommodates the prior text, [^2] I.e., establishing the accommodation exhibited by humans in the same setting. and the random baseline, which estimates the baseline level of accommodation in a dataset by replacing $r_{t}$ with $r_{rand}$, an utterance drawn from a random conversion and timestep in the same dataset.
@@ 3.1 Linguistic Indicators of Convergence
We measure the following features to characterize whether models alter their linguistic style to match that of their users, drawing on both human accommodation research and other common stylometric features for linguistic analysis Lagutina et al. (2019):
@@
LIWC Agreement A standard measure of linguistic accommodation is the frequency of LIWC (Linguistic Inquiry and Word Count) Chung and Pennebaker (2012) function word classes (e.g. Danescu-Niculescu-Mizil and Lee, 2011). Here, we consider the LIWC2007 classes considered in Ireland et al. (2011) (personal and impersonal pronouns, articles, conjunctions, prepositions, auxiliary verbs, frequently used adverbs, negations, and quantifiers), also using the $LSM$ metric to calculate how well each LIWC category distribution in generated responses aligns with prior turns. We both report the mean LIWC agreement across categories and provide a fine-grained analysis of each category in Section 5.3.
PROPN Overlap We calculate the overlap (percentage) of proper nouns between the text generated by the model and the preceding turn. We expect that language users who converge more with their interlocutor will have greater overlap in salient semantic topics, as approximated here by proper nouns.
Token Novelty We evaluate the percentage of tokens novel relative to the reference utterance; a lower percentage of novel tokens indicates that the model is converging more towards the user, or exhibiting greater lexical alignment Pickering and Garrod (2004). This is measured as $|w_{t}\cap w_{t-1}|/|w_{t}|$, where $w_{x}=\{w\in r_{x}\}$; prior work apply similar metrics of token novelty (or, inversely, token overlap) to measure lexical alignment between interlocutors (Ward and Litman, 2007; Duplessis et al., 2017, i.a.) .
@@ 4 Experimental Setup
Datasets We perform our dialogue prompting experiments on three popular English datasets: DailyDialog Li et al. (2017), containing conversations about daily life as written by English language learners; NPR Majumder et al. (2020), a dataset of radio interview transcripts; and the Movie Corpus Danescu-Niculescu-Mizil and Lee (2011), which contains conversations scraped from movie scripts. Each dataset covers a different dialogue domain; due to their respective data sources, all represent non-spontaneous dialogues (ranging from fully fictional conversations to constrained interviews).
@@ 5.3 Fine-grained Analysis of LIWC Categories
Table 2: Qualitative examples of model behavior on fine-grained stylometric features, drawn from DailyDialog experiments. Case A: Pretrained models overconverge on quantifiers; Case B: instruction-tuned models overuse (and thus underconverge on) conjunctions.
@@ 5.4 Stepwise Analysis of Linguistic Convergence
We find that human convergence scores fluctuate across timesteps; this is unsurprising, as alternating timesteps are uttered by the same speaker $S_{y}$ as the considered utterance $r_{t}$, indicating that each speaker’s linguistic patterns agree more with themselves than with their interlocutors. Interestingly, we see similar patterns across time with the language models for most metrics. This finding suggests that not only do language models adapt to their context, they also differentiate this adaptation to the particular speaker role Sacks et al. (1974) they are replacing within the dialogue. [^6] Interlocutors generally take different roles (speaker, listener) to coordinate turn-taking in conversation. While these models demonstrate sensitivity to speaker roles when emulating existing speakers, it remains an open question whether they maintain distinct speaker role patterns in open-ended dialogues without a specific conversational partner to emulate.
@@ 6 Discussion
Throughout our analysis, we compare model convergence to that of the human baselines to contextualize the models’ behavior. However, while in some cases the model converges to its context similarly to human utterances, we emphasize that the observed similarities do not necessarily indicate that the underlying causes of these behaviors are the same. Human accommodation is driven by speakers (often unconsciously) altering their speech to foster social and communicative goals Giles et al. (1991). In contrast, language models do not have these same underlying psychological communication goals when generating text.
@@
Instead, an important consideration is how the user will interpret texts from models that appear to accommodate them as a human interlocutor would, such as in the case of the “human-like" behavior we observe in instruction-tuned models. Appropriate model style has been shown to facilitate successful chatbot interactions Chaves et al. (2019); Thomas et al. (2020), and Bhatt and Rios (2021) finds that users tend to accommodate models more when they successfully produce topically relevant outputs, treating them more like human conversation partners. Other work has observed that while sycophantic behavior in langauge models can lead to lower model trust if detected by the user Carro (2024), users who don’t identify sycophancy in model generations will often overrely on these unhelpful model responses Bo et al. (2025). Thus, model convergence that better aligns with human behavior will likely mislead users to trust these models more, even if a model is unreliable for a given task in practice.
@@ 7 Conclusion
Thus, we consider the extent to which model convergence is even related to human accommodation. Our finer-grained analysis of LIWC categories (a prototypical feature for studying accommodation) shows that models exhibit very different patterns on these features than humans, suggesting that the underlying mechanisms for these behaviors are likely very different. Therefore, while this work primarily considers human behavior as a reference for characterizing model generations, future work should further characterize these differences and examine the underlying causes of the observed convergence in LMs.
@@
We approximate LLM participation in user-driven dialogues by having them complete turns in existing dialogue datasets. While this approach has experimental advantages (e.g., we can directly compare human accommodation features with the model’s responses in the same context), it also presents some limitations. Specifically, the model’s responses may be biased because it participates only in the later portion of (fixed) conversations; these conversational structures are also likely to differ significantly from those in natural interactions between a user and an LLM. Future work should supplement our experimental setting with more focused user studies to validate whether our findings hold in these cases. We also note that, though we perfrom our experiments on three popular, high-quality English dialogue datasets, each dataset’s source data imposes constraints on natural conversation structures, potentially limiting generalization of our findings.
While we perform experiments on sixteen language models, it remains an open question how larger models (>>70B parameters) and models post-trained on other objectives adapt to their users. The considered datasets cover various styles of conversations, but these differ from how some users interact with the models (i.e., information-seeking). Thus, future work should confirm whether these findings hold up in interactive user studies. Finally, while we test for an array of stylometric features, the model’s behavior of other aspects of style may differ.